OpenStudy (anonymous):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 25 y = 2x - 5 -4 and 0 4 and 0 -3 and -5 3 and -5

OpenStudy (anonymous):

@wio @eliassaab

OpenStudy (anonymous):

Use substitution. x^2 + (2x-5)^2 = 25

OpenStudy (anonymous):

Then solve for x.

OpenStudy (anonymous):

If you try to solve \[ x^2 + ( 2x -5)^2 =25 \] You find x=0 and y=4 x=0 gives y=5 x=4 gives y=3 These are not any of your choices unless I messed up somewhere

OpenStudy (anonymous):

@eliassaab ahh thanks i was confused with the substitution part can u help with this? A company sells boxes of duck calls (d) for $35 and boxes of turkey calls (t) for $45. They make batches of duck calls that fill 6 boxes and batches of turkey calls that fill 8 boxes. The company only has 42 boxes. Which system of equations helps the company plan to make $300? 6d = 42 - 8t 35d + 45t = 300 d = 42 - t 6d + 8t = 42 d = 42 - t 35d + 45t = 300 6d = 42 - 8t 35d - 45t = 300

OpenStudy (anonymous):

Correction You find x=0 and y=4 x=0 gives y=-5 x=4 gives y=3 These are not any of your choices unless I messed up somewhere

OpenStudy (anonymous):

Please post it as a new problem. It will avoid confusion with other posters.

OpenStudy (anonymous):

ok